The molarity of NaOH
1
Expert's answer
2020-05-24T03:10:53-0400

The molarity is defined as a number of moles of a certain substance in 1L of a solution. Molarity can be found by the following formula:

M = n / V = m / (Mr × V)

where n - number of moles, V - volume of a solution, m - mass, Mr - molecular mass.

The molecular weight of NaOH equals:

Mr(NaOH) = 23 g/mol + 16 g/mol + 1 g/mol = 40 g/mol

As a result, the molarity of NaOH can be calculated as following:

M(NaOH) = m / (40 g/mol × V).